OpenStudy (anonymous):

A camera manufacturer spends $1,800 each day for overhead expenses plus $9 per camera for labor and materials. The cameras sell for $18 each. a. How many cameras must the company sell in one day to equal its daily costs? b. If the manufacturer can increase production by 50 cameras per day, what would their daily profit be?

OpenStudy (anonymous):

let x be the number of cameras a. cost = 1800 + 9x revenue = 18x Solve for x: \(1800 + 9x = 18x\)

OpenStudy (anonymous):

200 now i need B

OpenStudy (anonymous):

for B, x = 250. profit =revenue - cost = \(18x-(1800+9x)\) plug in 250 for x